Solve for x.
x^( 2) + 8x + 16 = 25


Answer by deborabr(181)

x^2 + 8x + 16 = 25
x^2 + 8x + 16 -25 = 0
x^2 + 8x -9 = 0
(8)^2 -4*(1)*(-9)
64 + 36 = 100
-(8)+/- root of 100
___________________
2*(1)
-8 +/- 10
_________
2
x' = (-8-10)/2 = -18/2 = -9
x" = (-8+10)/2 = 2/2 = 1